The number x(x²+1),-1/2x²and 6 are the consecutive terms of an ap if x be real then the next two terms of the AP are
 

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
use property of 2b = a + c
 
now you can find the value of x and hence also value of d
 
hope it helps
Arun
5 years ago
Dear student
- x^2  = x^3 + x + 6
x^3 + x^2 + x+ 6 = 0
it is always increasing function hence only one real root is possible.
x = – 2 is only solution.
hence terms are
- 10  , – 2 , 6
common difference is 8.
hence next two terms are 14 and 22
 
Regards
Arun (askIITians forum expert)